OpenStudy (onepieceftw):

Find all the solutions within interval [0, 2pi) 2 sin^2 x = sin^2 x

OpenStudy (mrhirohito):

@mathmale

Nnesha (nnesha):

Set it equal to zero. Solve for sin x.

OpenStudy (onepieceftw):

2 sin^2x - sinx = 0 I basically clueless at this stage.

OpenStudy (onepieceftw):

correction: 2 sin^2x - sinx^2 = 0

OpenStudy (onepieceftw):

sin^2x

Nnesha (nnesha):

that's correct lets say sin x =y so if we replace sinx with y \[\large\rm 2y^2-y^2=0\] how would you solve this equation ?

OpenStudy (onepieceftw):

Wouldn't we take the square root of both sides before setting it equal to 0?

Nnesha (nnesha):

no. first we need to set it equal to zero. now as you can the variables of both terms are the same \[\large\rm \rm 2\color{red}{y^2}-1\color{Red}{y^2}=0\] combine like terms ? ye?

Nnesha (nnesha):

see*

OpenStudy (onepieceftw):

Oh okay. I had it written down on my paper wrong so i didn't see them as like terms.

OpenStudy (onepieceftw):

so sin^2x = 0

Nnesha (nnesha):

yep

OpenStudy (onepieceftw):

Where does it go from there?

Nnesha (nnesha):

how would you get rid of the square(^2)) ???

OpenStudy (onepieceftw):

Take the square root but there's nothing on the right side of the equation

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @OnePieceFTW so sin^2x = 0 \(\color{blue}{\text{End of Quote}}\) ^^^^ there is 0 at right side.

Nnesha (nnesha):

and yes take square root both sides

OpenStudy (anonymous):

\[2 \sin ^2x-\sin ^2x=0,~or~\sin ^2x=0,2\sin ^2x=2*0=0,1-\cos 2x=0, \cos 2x=1\] \[\cos 2x=1=\cos 0,\cos 2 \pi,\cos 4 \pi,2x=0,2 \pi, 4 \pi,x=0,\pi,2 \pi\]